OpenStudy (anonymous):

2. Explain how the order of operations determines how you simplify a numeric expression. Be sure to explain the rules for addition, subtraction, multiplication, division, and the use of grouping symbols. Use examples to illustrate your points.

OpenStudy (anonymous):

If anyone has to do this in the future I finished it myself and if you want to use it change the words a bit or if its not useful dont use it and figure it out yourself (its pretty simple but im pretty lazy haha) here it is, its just a short explanation with some examples on how its use Order of operations is a rule to solve a problem, to solve a order of operation problem you will use PEMDAS which stands for Parentheses, Exponent, Multiplication, Division, Addition, Subtraction. You use it in order from 'P' til 'S' if the problem contains it if it does not then skip to the next letter. A given problem such as 9 x (5 – 2) is very simple with PEMDAS. You begin with 'P' which stands for parentheses. ( 5 – 2) = 3, after you solve that you still have 9 x 3 and with this you simply multiply it, 9 x 3 = 27. Another one would be 21 – 3^2 x (8 – 6), you start with parentheses (8 – 6) = 2, solve the next which would be the exponent 3^2 = 3 x 3 = 9, so far it is 21 – 9 x 2, now we have to multiply 9 x 2 = 18, and 21 – 18 = 3 which is our solution.